Two lightbulbs are 2.0 m apart. Part A From what distance can these lightbulbs be marginally resolved by a small telescope with a 3.0-cm -diameter objective lens? Assume that the lens is limited only by diffraction and λ = 600 nm.
Answer #1

λ = 600x10^-9 m

Raleigh’s' criteria for angular resolution of a circular aperture of diameter,

D>>λ, is:

(dθ) = 1.22λ/D = 1.22*600x10^-9/0.03 = 24.4x10^-6

Also, (dθ) = d/L,

=> L = 2/24.4x10^-6 = 8.2x10^4 m (= 82 km)

Kindly rate :)